40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue May 1 03:12:58 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_0849.jpg
    Reply  Tue May 1 23:58:21 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_0851.jpg
       Reply  Wed May 2 19:48:32 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_0853.jpgIMG_0854.JPGIMG_0852.jpgresult.png
          Reply  Thu May 3 22:26:14 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_0855.JPGIMG_0856.jpg
             Reply  Fri May 4 20:19:58 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_0863.JPGIMG_0864.jpg
                Reply  Wed May 9 03:11:33 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IO_lid_compare.pngIO_lid_compare.pngcomputer_compare.png
                   Reply  Wed May 30 23:39:15 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_1300.jpg
                   Reply  Wed May 30 23:39:41 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam IMG_1300.jpg
                      Reply  Fri Jun 1 01:00:30 2012, frank, tara, DailyProgress, Acoustics, acoustic enclosure for input beam beat_2012_05_31.pngbeat_2012_05_31_b.png
Message ID: 975     Entry time: Fri Jun 1 01:00:30 2012     In reply to: 973
Author: frank, tara 
Type: DailyProgress 
Category: Acoustics 
Subject: acoustic enclosure for input beam 

The beat signal is measured with an adhoc lid covering the enclosure on the beat part. There are no significant improvement in the acoustic region of the beat signal yet.

     The closing lid is the only missing piece for our acoustic box, so I just use smaller pieces of acrylic planes I can find to cover the top, and measure the beat signal. There is not much improvement in the acoustic region, see fig2 below for a closer look.

     There might  be dust on the optics somewhere, because the signal at low frequency is not as stable as it used to be. I could see scattered light bump coming up when I averaged for too long (this did not happen last time). I'll check all optics and make sure that they are clean.

     NOTE: The temperature servo is back on. We had to disconnect it in order to install the acoustic box. I set C3:PSL-RCAV_TEMPAVG to 32.5 ( the previous value was 35.03). I changed it because I noticed that as the temperature reached up to the set point, the beat frequency went up as well.  I'll check if this set point will reduce beat frequency down to ~ 160 MHz or not.

    

 

beat_2012_05_31.png

beat_2012_05_31_b.png

ELOG V3.1.3-